OpenStudy (anonymous):

If 8 sandwich rolls cost $2.25, how much will 21 rolls cost? Please show all work.

OpenStudy (anonymous):

Ok, so if 8 sandwich cost 2.25 how would you find out what one sandwich cost?

OpenStudy (anonymous):

Well 8 divided by 2.25 is 3.55 so each one is 3.55 so 3.55 times 21 = 74.66 or rounded $74.70

OpenStudy (anonymous):

8 divided by $2.25

OpenStudy (anonymous):

Good. Now you know How much each sandwich cost. To figure out how much 21 sandwich cost multiply by 21

OpenStudy (anonymous):

Ummm...no lol. 2.25/8 will tell you how much each costs. It wouldn't make sense to have 8 for 2.25 but they are 3.55 EACH?!? Lol. So you have 2.25/8=0.28*21=5.88

OpenStudy (anonymous):

thanks. much appreciated.